A bar has a sign saying all mixed drinks contain 1.25 oz of alcohol. Using the equivalences 16 oz = 1 lb ; 1 lb = 454 g ; 0.789 g alcohol = 1 mL, what is the volume of the drink in mL if the drink is 11.2 % alcohol by volume (X mL alcohol = 100 mL drink)?
1
Expert's answer
2018-01-18T13:48:07-0500
1 oz = 1/16 lb = 0.0625 lb = 28.375 g
1.25 oz = 35.47 g
Volume of alcohol: 35.47 / 0.789 = 44.96 mL
Volume of the drink: 44.96 / 0.112 = 401.4 mL